The larger of two numbers is 8 more than four times the smaller. If the larger is increased by four times the smaller, the result is 40. Find the numbers.
Answer by usyim88hk(158)

Let x be the smaller number:
x = smaller number
4x+8 = bigger number
4x+8+4x = 40 (If the larger is increased by four times the smaller, the result is 40)
8x+8 = 40
8x = 32
x = 4
So the small number is 4 and the bigger number is 4(4)+8 = 24